OpenStudy (anonymous):

Not completely sure about the answer. Can someone help solve? -16t2 + 64t + 4

OpenStudy (anonymous):

If you have to factor it the answers -4(4t^2-16t-1)

OpenStudy (anonymous):

I need find a solution by either using the quadratic formula, completing the square or factoring in completely.

OpenStudy (anonymous):

That's factoring it completely

OpenStudy (phi):

I assume you are solving -16t2 + 64t + 4 = 0 ? as pointed out above, you can factor a -4 out from each term to get \[ -4(4t^2 -16t -1)=0 \] divide both sides by 4 to get \[ 4t^2 -16t -1)=0 \] now use the quadratic formula to find the two values of "t" that make this equation true. See http://www.khanacademy.org/math/trigonometry/polynomial_and_rational/quad_formula_tutorial/v/using-the-quadratic-formula for how to do this

OpenStudy (anonymous):

plz help

OpenStudy (nikato):

do u know what the quadratic formula is?